the sides of a quadrilateral are consecutive numbers. if the perimeter of the quadrilateral is 146 cm, find the lenths of the sides
Answer by oberobic(2304)

4 consecutive numbers can be defined as:
x
x+1
x+2
x+3
------
4x + 6 = 146
4x = 140
x = 35
.
x+1 = 36
x+2 = 37
x+3 = 38
.
Answer: The sides have length: 35, 36, 37, and 38 cm.